Jen is 80 years old and her grandson mark is 5 years old, in how many years will Jen be 4 times Mark's age ?
Answer by macston(5194)

.
J=Jen's age=80 years; M=Mark's age=5 years;
x=years until Jen is 4 times Mark's age
.
J+x=4(M+x)
80+x=4(5+x)
80+x=20+4x
60=3x
20=x
ANSWER: Jen will be 4 times Mark's age in 20 years.
.
CHECK:
J+x=4(M+x)
80+20=4(5+20)
100=4(25)
100=100